I often have trouble with fractions and the correct method to solve. I would like to have work shown to match my answers and see if they are correct
1)Find the difference between 18 1/2 and 14 2/3? (3 5/12 =answer)
2) Jim walked 3/8 mile to the grocery store 5/6 mile from there to the post office 1/3 mile back home. How far did he walk? (37/24 miles)
3) John had 2/3 gallon of milk used 3/8 of it. How much was used? (1/4 gallon)
4) Average of 1/4 and 1/2= (3/8)
5) find the product of 3/5 and 2/3= 1/3

Answer by Alan3354(69443)

1)Find the difference between 18 1/2 and 14 2/3? (3 5/12 =answer)
= 18 3/6 - 14 4/6
= 17 9/6 - 14 4/6
= 3 5/6
-----------------
2) Jim walked 3/8 mile to the grocery store 5/6 mile from there to the post office 1/3 mile back home. How far did he walk? (37/24 miles)
3/8 + 5/6 + 1/3
= 9/24 + 20/24 + 8/24
= 37/24 or 1 13/24
--------------------
3) John had 2/3 gallon of milk used 3/8 of it. How much was used? (1/4 gallon)
(2/3)*(3/8) = 6/24 = 1/4
---------------------
4) Average of 1/4 and 1/2= (3/8)
correct
---------------
5) find the product of 3/5 and 2/3= 1/3
= 2/5